# Lumina-mGPT: Illuminate Flexible Photorealistic Text-to-Image Generation with Multimodal Generative Pretraining

### v1

@@ Abstract
We present Lumina-mGPT, a family of multimodal autoregressive models capable of various vision and language tasks, particularly excelling in generating flexible photorealistic images from text descriptions. Unlike existing autoregressive image generation approaches, Lumina-mGPT employs a pretrained decoder-only transformer as a unified framework for modeling multimodal token sequences. Our key insight is that a simple decoder-only transformer with multimodal Generative PreTraining (mGPT), utilizing the next-token prediction objective on massive interleaved text-image sequences, can learn broad and general multimodal capabilities, thereby illuminating photorealistic text-to-image generation. Building on these pretrained models, we propose Flexible Progressive Supervised Finetuning (FP-SFT) on high-quality image-text pairs to fully unlock their potential for high-aesthetic image synthesis at any resolution while maintaining their general multimodal capabilities. Furthermore, we introduce Ominiponent Supervised Finetuning (Omni-SFT), transforming Lumina-mGPT into a foundation model that seamlessly achieves omnipotent task unification. The resulting model demonstrates versatile multimodal capabilities, including visual generation tasks like flexible text-to-image generation and controllable generation, visual recognition tasks like segmentation and depth estimation, and vision-language tasks like multi-turn visual question answering. Additionally, we analyze the differences and similarities between diffusion-based and autoregressive methods in a direct comparison. All code and checkpoints will be available on our GitHub repository: https://github.com/Alpha-VLLM/Lumina-mGPT.
@@ 1 Introduction
Seminal models, including DALL-E 3 (Betker et al., 2023), Stable Diffusion 3 (Esser et al., 2024), and SoRA (Brooks et al., 2024), have demonstrated superior performance in photorealistic image and video generation using diffusion-based generative modeling over continuous latent image features. In contrast, autoregressive (AR) generative models, which rely on “next-token prediction,” have revolutionized text generation with groundbreaking reasoning abilities, as exemplified by models like GPT-4 (Achiam et al., 2023) and Gemini (Team et al., 2023), over discrete token representation. However, AR-based generative modeling over vector-quantized image features still lags far behind diffusion-based counterparts in terms of photorealistic and controllable image generation.
Although previous autoregressive efforts, such as DALL-E (Ramesh et al., 2021), CogView (Ding et al., 2021), Parti (Yu et al., 2022), OFA (Wang et al., 2022), Unified-IO (Lu et al., 2022; 2024a), LlamaGen (Sun et al., 2024), and Chameleon (Team, 2024a), have extensively explored generative modeling over vector-quantized image features following the paradigm of large language models (LLMs), their results on text-to-image generation have either been unsatisfactory or produced high-quality samples limited to academic benchmarks like ImageNet (Deng et al., 2009). We summarize the limitations of previous approaches below:
Randomly-Initialized Transformer While transfer learning has revolutionized key fields such as visual recognition (He et al., 2016; Radford et al., 2021; Lu et al., 2019) and language generation (Raffel et al., 2020; Radford et al., 2018; 2019; Brown et al., 2020), popular autoregressive image generation approaches such as DALL-E, Parti, and LlamaGen all adopt a randomly-initialized causal transformer, which fails to utilize pretrained transferable representation and large-scale datasets. As a result, AR-based approaches often lead to poor image generation quality and slow convergence without leveraging proper large-scale pretraining.
Verbose Encoder-Decoder Architeture DALL-E and CogView initially propose using a decoder-only transformer for image generation with discrete representation, where a single transformer acts as both a text encoder and an image token decoder. However, subsequent approaches, such as Parti and LlamaGen, adopt a verbose encoder-decoder architecture that injects frozen T5 text features (Raffel et al., 2020) using cross-attention or prefix-filling approaches, motivated by the findings of Imagen (Saharia et al., 2022). Compared to the trend in scaling LLMs (Touvron et al., 2023; Bai et al., 2023; Bi et al., 2024), such encoder-decoder architecture is cumbersome due to the decoupling of text encoding and image token modeling. This design significantly complicates the autoregressive-decoding framework, limits the scalability of image generation, and hinders the generalization to additional modalities and tasks.
Limited Decoding Resolution and Flexibility Natural images exsit in various resolutions and aspect ratios. Advanced diffusion models (Chen et al., 2024a; b; Esser et al., 2024; Team, 2024b; Li et al., 2024b; Zhuo et al., 2024; Lu et al., 2024b) can successfully generate diverse photorealistic images at arbitrary resolution with skewed ratios. In contrast, current AR-based approaches (Yu et al., 2022; Team, 2024a; Sun et al., 2024) rely on central-cropping a low-resolution 512 $\times$ 512 image and transforming the cropped low-resolution image into a fixed-length sequence of discrete tokens using a pretrained Vector-Quantized Variational Autoencoder (VQ-VAE) (Van Den Oord et al., 2017; Razavi et al., 2019; Esser et al., 2021). This approach simplifies autoregressive training but at the cost of deteriorated image quality and generation flexibility.
Poor Task Extensability Autoregressive modeling excels at unified generative modeling for various tasks and modalities using discrete tokens (Wang et al., 2022; Lu et al., 2022; 2024a). However, previous AR-based image generation approaches (Ramesh et al., 2021; Ding et al., 2021; Yu et al., 2022; Sun et al., 2024) have been limited to text-to-image generation without exploring the unification with other tasks, such as dense labeling and controllable image generation. This lack of task extensibility significantly constrains the applicability of AR-based image generation models to a broader range of scenarios.
@@
Table 1: Overview of the design choices and capabilities of multimodal autoregressive approaches. Lumina-mGPT is the only model capable of both flexible photorealistic image generation and multimodal task unification, due to its decoder-only transformer design and multimodal generative pretraining.
Method Model Multimodal Flexible Photorealistic Task
Architecture Pretraining Image Generation Extensibility
DALL-E (Ramesh et al., 2021) Decoder-only × × ×
Cogview (Ding et al., 2021) Decoder-only × × ×
Parti (Yu et al., 2022) Encoder-Decoder × × ×
LlamaGen (Sun et al., 2024) Encoder-Decoder × × ×
OFA (Wang et al., 2022) Encoder-Decoder ×
Unified-IO (Lu et al., 2022) Encoder-Decoder × ×
Unified-IO 2 (Lu et al., 2024a) Encoder-Decoder ×
Chameleon (Team, 2024a) Decoder-only × ×
@@
To address the aforementioned challenges, we present Lumina-mGPT, a decoder-only transformer initiated with effective multimodal Generative PreTraining (mGPT) and then supervised-finetuned over flexible, high-quality, high-resolution discrete image tokens in a progressive manner. This framework illuminates flexible high-resolution photorealistic image generation and can be easily extended to solve various component tasks in a unified manner. We provide a detailed comparison of the architecture design choices and model capabilities of existing multimodal autoregressive approaches in Table 1. The key features of Lumina-mGPT are outlined below:
@@
➀ Effective Multimodal Generative Pretraining Unlike commonly adopted approaches that utilize a randomly-initialized causal transformer to generate discrete image tokens in an autoregressive manner, Lumina-mGPT is initialized from an effective multimodal generative pretraining representation. mGPT is a multimodal autoregressive transformer trained on a large-scale image-text interleaved corpus using a simple “next-token prediction” loss. Our key finding is that leveraging the rich knowledge from large-scale multimodal pretraining, mGPT can learn broad and general multimodal representation. This serves as an excellent starting point, remarkably simplifying the task of text-to-image generation, accelerating training convergence on downstream tasks, and unifying various vision-language tasks. Due to the lack of training resources, this mGPT representation is directly adopted from the pretrained Chameleon 7B and 30B models released by Meta (Team, 2024a).
➁ Simple Decoder-Only Architecture As mGPT is pretrained on multimodal tokens, it naturally learns vision and language representation within a single transformer. Consequently, with one decoder-only causal transformer, Lumina-mGPT can handle both text encoding and image token decoding simultaneously. By sharing this decoder-only framework with text-only LLMs, Lumina-mGPT can benefit from well-established theories and infrastructures in the LLM community, such as scaling properties (Chowdhery et al., 2022; Brown et al., 2020) and techniques to optimize the training and inference processes (Dao et al., 2022; Dao, 2024; Kwon et al., 2023). We train a family of Lumina-mGPTs with up to 30 billion parameters to provide a flexible trade-off between efficiency and performance and to explore the limit of its multimodal capabilities as model parameters scale.
➂ Flexible High-Resolution Decoding Benefiting from large-scale pretraining on multi-modal interleaved datasets and advanced architectural designs including query-key normalization (QK-Norm) (Henry et al., 2020; Lu et al., 2024a; Zhuo et al., 2024; Team, 2024a) and rotary positional embedding (RoPE) (Su et al., 2024), the mGPTs are ideal initializations for flexible photorealistic image generation. We thus propose Flexible Progressive Supervised Finetuning (FP-SFT) to fully exploit the potential of mGPT for flexible, photorealistic high-resolution text-to-image generation. This approach starts with low-resolution discrete tokens and progressively transitions to high-resolution discrete tokens. Combined with resolution-aware prompts, this weak-to-strong SFT strategy effectively and efficiently enhances the image generation capabilities of Lumina-mGPT. To equip Lumina-GPT with the ability to understand and generate images at flexible resolution and aspect ratio, we also propose Unambiguous image Representation (Uni-Rep), which eliminates the ambiguity in the original 2D shape of the flattened 1D image tokens by introducing height/width indicators and <end-of-line> tokens.
➃ Omnipotent Task Unification One advantage of AR-based approaches over diffusion-based generative modeling is their promising potential for unified sequence modeling across various modalities and tasks within a single framework (Chen et al., 2022; Wang et al., 2022; Lu et al., 2022; 2024a). In this paper, we explore this potential through preliminary experiments. Specifically, we introduce Omnipotent Supervised Finetuning (Omni-SFT), which treats various tasks–such as multi-turn dialog (Achiam et al., 2023; Meta, 2024; Bai et al., 2023; Bi et al., 2024), visual multi-turn understanding (Liu et al., 2023; Chen et al., 2024c), dense labeling (Carion et al., 2020; Kirillov et al., 2023; Yang et al., 2024), flexible image generation (Zhuo et al., 2024; Ren et al., 2024; Du et al., 2024), and spatial-conditional image generation (Zhang et al., 2023; Ye et al., 2023)–as a unified discrete modeling task. We then perform Omni-SFT over mGPT to seamlessly achieve omnipotent task unification via a natural language interface.
We summarize our contributions as follows: (1) We discover that autoregressive transformers with multimodal generative pertaining, rather than being trained from scratch, can serve as an effective initialization for photoelastic text-to-image generation and the unification of vision and language tasks. (2) Building on mGPTs, we introduce two novel finetuning strategies, FP-SFT and Omni-SFT, to fully unleash their potential in a weak-to-strong paradigm using only 10M high-quality text-image data. (3) The resulting model, Lumina-mGPT, not only demonstrates impressive photorealistic text-to-image generation at any resolution, bridging the gap between autoregressive and diffusion-based methods, but also seamlessly supports various downstream tasks using a language interface.
2 Related Work
Multimodal Large Language Models
Recent large language models (LLMs) (Chowdhery et al., 2022; Touvron et al., 2023; Brown et al., 2020; Achiam et al., 2023; Team et al., 2023) demonstrate strong instruction-following and reasoning capabilities, coupled with extensive world knowledge. To extend LLMs’ expertise from the text domain to multimodal domains such as images and videos, prior works (Liu et al., 2023; Lin et al., 2023a; Maaz et al., 2023; Lin et al., 2023b) have aligned pretrained encoders from various modalities with LLMs by curating multimodal instruction tuning datasets. Although these multimodal large language models (MLLMs) have exhibited powerful visual understanding capabilities, their multimodal functionality is primarily limited to perceiving the visual world, exemplified by tasks such as visual question answering and image captioning, rather than generating visual outputs. Another line of research Koh et al. (2024); Dong et al. (2024); Sun et al. (2023); Wu et al. has proposed augmenting MLLMs with the ability to generate images, videos, and audio from text instructions. These approaches introduce additional visual tokens for generation and align these generative tokens as conditional information with a pretrained generator, such as Stable Diffusion (Esser et al., 2024; Podell et al., 2023) for text-to-image generation. Consequently, the generation capabilities heavily rely on the external expert generator rather than MLLMs themselves, resulting in inconsistent and inferior generation results. To combine the strength of both approaches, our model aims to learn both understanding and generation of images using an MLLM with native multimodal capabilities, drawing inspiration from Chameleon (Team, 2024a), a mixed-modal early-fusion foundation model.
Text-to-Image Generation
The task of text-to-image generation seeks to synthesize photorealistic and diverse images based on textual descriptions. Nowadays, diffusion models, whether in pixel space (Saharia et al., 2022) or in latent space (Podell et al., 2023; Esser et al., 2024; Pernias et al., 2024), have become the de-facto approaches due to their superior performance, particularly in producing extremely high-aesthetic images. Among these models, the recent trend of scaling diffusion transformers (DiTs) (Chen et al., 2024b; a; Gao et al., 2024; Esser et al., 2024; Li et al., 2024b; Team, 2024b) suggests a unified architecture for both text and image modeling. However, existing DiTs still leverage separate language models, such as CLIP (Radford et al., 2021) or T5 (Raffel et al., 2020), as text encoders. This modality gap between text and image representations not only leads to inaccurate generation but also hinders the development of a unified multimodal foundational generative model. Compared to the dominance of diffusion models, the progress of autoregressive image generation has received less attention in recent years. Early works (Ramesh et al., 2021; Ding et al., 2021) proposed a two-stage generation approach: first, training a VQ-VAE Van Den Oord et al. (2017); Esser et al. (2021) for image tokenization and de-tokenization, and then using an autoregressive transformer to model discrete image token sequences, akin to language modeling. Parti (Yu et al., 2022) scaled up the autoregressive transformer to 20 billion parameters, demonstrating promising high-fidelity image generation results. LlamaGen (Sun et al., 2024) further improved the image tokenizer and integrated advanced techniques in LLMs, bridging the performance gap with diffusion counterparts. Unlike Parti and LlamaGen, Lumina-mGPT proposes multimodal generative pertaining on unified text-image sequences, followed by supervised finetuning on high-quality text-to-image pairs, achieving flexible high-aesthetic image generation with autoregressive models.
@@ 3 Methodolgy
Lumina-mGPT is a decoder-only transformer initialized with multimodal Generative PreTraining (mGPT) and finetuned over high-quality multimodal tokens derived from various tasks. Based on the robust mGPT representation and our proposed supervised finetuning strategies with unambiguous image representation, Lumina-mGPT achieves superior performance in photorealistic image generation and omnipotent task unification with high flexibility in image resolution and aspect ratio. In this section, we first introduce mGPT, followed by the training and inference details of Lumina-mGPT.
@@ 3.1 Revisiting mGPT with Chameleon
mGPT represents the family of models utilizing a decoder-only transformer architecture, pretrained on extensive multimodal token sequences. These models exhibit exceptional native multimodal understanding, generation, and reasoning capabilities, offering the potential for universal modeling across various modalities and tasks. We use the recent open-source model, Chameleon (Team, 2024a), as an example to illustrate the design choices and implementing details of mGPT.
@@
To unify text and images into a multimodal token sequence, it is essential first to tokenize both text and image into discrete space. Especially for images, the choice of tokenizer is crucial as it determines the upper limit of generation quality. Specifically, Chameleon trains a new byte pair encoding tokenizer for text. For images, it adopts the quantization-based tokenization method following prior works (Van Den Oord et al., 2017; Razavi et al., 2019; Esser et al., 2021; Yu et al., 2022; Ramesh et al., 2021), converting continuous image patches into discrete tokens from a fixed codebook while reducing spatial dimensions. The quantized image tokens are then flattened into a 1D sequence and concatenated with text tokens in various ways to form a multimodal token sequence for unified modeling.
@@
Unlike Unified-IO and Parti using an encoder-decoder architecture with pretrained encoders, mGPT trains a decoder-only autoregressive transformer from scratch after transforming text and image inputs into a unified sequence of discrete tokens $x=(x_{1},x_{2},...,x_{T})$, leading to a simpler and more unified approach for multimodal generative modeling. mGPT adopts a standard dense transformer architecture for scalability and generalizability, with minor adaptations like RoPE (Su et al., 2024) and SwiGLU (Shazeer, 2020) activation function, following the LLaMA series (Touvron et al., 2023).
However, this standard transformer architecture exhibits training instability when scaling up the model size and context length of multimodal token sequences, as observed in both Chameleon (Team, 2024a) and Lumina-Next (Zhuo et al., 2024). It has been found that this instability is caused by the uncontrollable growth of network activations in transformer blocks. Therefore, Pre-Norm, Post-Norm, and QK-Norm are added to each transformer block to preserve the magnitude of intermediate activation and further stabilize the training process.
Training Objective
During training, mGPT models the conditional probability $p(x_{t}|x_{1},...,x_{t-1})$ of multimodal sequences using the standard next-token prediction objective. Additionally, Chameleon applies z-loss (Chowdhery et al., 2022) to stabilize the training of 7B and 30B models. Initially, we underestimated the importance of z-loss in our attempts because it is absent in most works relevant to the training of (multimodal) LLMs. However, we found that without this term, the magnitude of logits surges, causing the loss to diverge. On the other hand, with z-loss, we observe that for the 7B and 30B models, the optimal temperature for inference-time image generation is much lower than that for the 1B model, as the magnitude of logits is significantly reduced in the larger models.
@@
Although mGPTs like Chameleon demonstrate potential for joint image and text understanding within one decoder-only transformer, their image generation ability remains inferior to state-of-the-art diffusion-based frameworks (Esser et al., 2024; Chen et al., 2024a; Li et al., 2024b; Team, 2024b; Zhuo et al., 2024) in both quality and resolution flexibility. Moreover, it is worth noting that the image generation ability is even absent in the open-source version of Chameleon. Additionally, the capabilities of Chameleon are confined to vision-language and text-only tasks, excluding a broader range of vision-centric tasks. These include classic visual recognition tasks such as segmentation and depth prediction, as well as creative visual generation tasks like controllable generation and image editing. Lumina-mGPT is built upon Chameleon to unlock its full potential for flexible photorealistic image generation and to become a versatile vision generalist.

### v3

@@ Abstract
We present Lumina-mGPT, a family of multimodal autoregressive models capable of various vision and language tasks, particularly excelling in generating flexible photorealistic images from text descriptions. By initializing from multimodal Generative PreTraining (mGPT), we demonstrate that decoder-only Autoregressive (AR) model can achieve image generation performance comparable to modern diffusion model with high efficiency through Flexible Progressive Supervised Finetuning (FP-SFT). Equipped with our proposed Unambiguous image Representation (Uni-Rep), Lumina-mGPT can flexibly generate high-quality images of varying aspect ratios. Building on the strong image generation capabilities, we further explore Ominiponent Supervised Finetuning (Omni-SFT), an initial attempt to elevate Lumina-mGPT into a unified multi-modal generalist. The resulting model demonstrates versatile multimodal capabilities, including visual generation tasks like text-to-image/multiview generation and controllable generation, visual recognition tasks like segmentation and depth estimation, and vision-language tasks like multi-turn visual question answering, showing the rosy potential of the technical direction. Codes and checkpoints are available at https://github.com/Alpha-VLLM/Lumina-mGPT.
Machine Learning, ICML
Figure 1: Unified in the next-token prediction framework, Lumina-mGPT can perform a wide range of multi-modal tasks. See Figure 6 to Figure 11 in the Appendix for more demonstrations.
@@ 1 Introduction
Seminal models, including DALL-E 3 (dalle3), Stable Diffusion 3 (sd3), and SoRA (sora), have demonstrated superior performance in photorealistic image and video generation using diffusion-based generative modeling over continuous latent image features. In contrast, autoregressive (AR) generative models, which rely on “next-token prediction”, have revolutionized text generation with groundbreaking reasoning abilities, as exemplified by models like GPT-4 (gpt4) and Deepseek-v3 (gemini).
However, AR-based generative modeling with vector-quantized image features remains a relatively less mature topic. While previous autoregressive efforts, such as DALL-E (dalle), CogView (cogview), Parti (parti), OFA (ofa), Unified-IO (unified-io; unified-io2), LlamaGen (llamagen), and Chameleon (chameleon), have made important explorations, each of these methods suffers from one or more of the following limitations: (1) their image generation capabilities have been either unsatisfactory or constrained to academic benchmarks such as ImageNet (imagenet); (2) they rely on complex architectures, such as encoder-decoder frameworks, which impede scalability and limit generalization to other tasks; (3) they are restricted in decoding resolution and flexibility, often producing small images with fixed resolutions; and (4) they lack task extensibility, making them unsuitable for broader scenarios that require integration with tasks like dense labeling or controllable image generation. A more detailed discussion of these challenges is provided in Appendix D.
@@
Table 1: Design choices and capabilities of multimodal autoregressive approaches. Lumina-mGPT is the only model capable of both flexible photorealistic image generation and multimodal task unification, due to its decoder-only transformer design and multimodal generative pretraining.
Method Model Multimodal Photorealistic Flexible Image Task
Architecture Pretraining Image Generation Aspect Ratio Extensibility
DALL-E (dalle) Decoder-only × × × ×
Cogview (cogview) Decoder-only × × × ×
Parti (parti) Encoder-Decoder × × × ×
LlamaGen (llamagen) Encoder-Decoder × × × ×
OFA (ofa) Encoder-Decoder × ×
Unified-IO (unified-io) Encoder-Decoder × × ×
Unified-IO 2 (unified-io2) Encoder-Decoder × ×
CM3Leon (cm3leon) Decoder-only × ×
Chameleon (chameleon) Decoder-only × × ×
@@
To address the aforementioned challenges, we present Lumina-mGPT, a decoder-only transformer initiated with effective multimodal Generative PreTraining (mGPT) and then supervised-finetuned over flexible, high-quality, high-resolution discrete image tokens in a progressive manner. This framework illuminates flexible high-resolution photorealistic image generation and can be easily extended to solve various downstream tasks in a unified manner. We provide a detailed comparison of the architecture design and model capabilities of existing multimodal autoregressive approaches in Table 1. The key features of Lumina-mGPT are outlined below:
➀ Simple Decoder-Only Architecture: We adopt a simple decoder-only architecture, which offers a significant advantage over more complex designs such as encoder-decoder architectures. Decoder-only models provide an elegant and extensible framework that unifies various understanding and generation tasks across different modalities. This makes them a compelling choice for achieving true unification. Furthermore, by leveraging the same architecture as the rapidly evolving text-only LLMs, Lumina-mGPT can take advantage of well-established theories and infrastructure in the LLM community, including scaling properties (palm; gpt3) and advanced techniques for optimizing training and inference (flashattention; flashattention2; vllm).
➁ Flexible High-quality Image Generation: Despite the aforementioned strengths of decoder-only architecture, the image generation capabilities of such models remain limited, creating a gap between the potential and the reality of this architecture. We thus propose Flexible Progressive Supervised Finetuning (FP-SFT) to fully fulfill the potential of high-quality text-to-image generation. This approach starts with low-resolution discrete tokens and progressively transitions to high-resolution discrete tokens. Combined with Unambiguous image Representation (Uni-Rep), this weak-to-strong SFT strategy effectively grants the model with the ability to generate high-quality photorealistic images with flexible aspect ratios.
➂ Omnipotent Task Unification: The high-quality image generation capabilities achieved during the FP-SFT stage provide the prerequisites for further exploring the fundamental strength, namely task unification, of the decoder-only architecture. We thus propose Omnipotent Supervised Finetuning (Omni-SFT), a preliminary attempt to create a multi-modal omnipotent generalist. Omni-SFT treats various tasks–such as multi-turn dialog, visual multi-turn understanding, dense labeling, text-to-image generation, text-to-multiview generation, image editing, and spatial-conditional image generation–as a unified discrete modeling task, allowing the model to achieve omnipotence via a unified natural language interface.
@@
➃ Effective Multimodal Generative Pretraining: Instead of random initialization, Lumina-mGPT leverages an effective multimodal Generative PretTraining (mGPT). This representation is derived from a multimodal transformer trained at scale using a straightforward "next-token prediction" loss. The rich knowledge encapsulated in this pretraining significantly accelerates the learning process for text-to-image generation and the downstream vision-language tasks. Given the constraints in training resources, we adopt the mGPT representation directly from the pretrained Chameleon 7B and 30B models released by Meta (chameleon).
The summarized contributions are as follows: (1) We are the first, particularly in the open-source domain, to demonstrate that a decoder-only AR model can achieve image generation performance on par with modern diffusion models. Besides, with mGPT initialization, this capability can be achieved with remarkable efficiency, requiring only 32 A100 GPUs over 7 days to train a 7B model. (2) We introduce UniRep, a novel image representation that enables decoder-only AR models to generate images with varying aspect ratios, offering greater flexibility in image generation. (3) Leveraging the strong image generation capabilities achieved, we propose Omni-SFT, a pioneering exploration aimed at elevating the model into a unified generalist. (4) We release the full-pipeline code implementation to facilitate further advancements from the research community.
@@ 2 Methodolgy
Lumina-mGPT is a decoder-only transformer initialized with multimodal Generative PreTraining (mGPT) and finetuned over high-quality multimodal tokens derived from various tasks. Based on the robust mGPT representation and our proposed supervised finetuning strategies with unambiguous image representation, Lumina-mGPT achieves superior performance in photorealistic image generation and omnipotent task unification with high flexibility in image resolution and aspect ratio.
@@ 2.1 Revisiting mGPT with Chameleon
mGPT represents the family of models utilizing a decoder-only transformer architecture and pretrained on extensive multimodal token sequences. We use the recent open-source model, Chameleon (chameleon), as an example to illustrate the design choices and implementing details of mGPT.
@@
To unify text and images into a multimodal token sequence, it is essential first to tokenize both text and image into discrete space. The choice of image tokenizer is especially crucial as it determines the upper limit of generation quality. Chameleon trains a byte pair encoding (BPE) tokenizer for text; for images, it adopts the quantization-based tokenization method following prior works (vqgan; parti; dalle), converting continuous image patches into discrete tokens from a fixed codebook while reducing spatial dimensions. The quantized image tokens are then flattened into 1D sequence and concatenated with text tokens to form a multimodal token sequence for unified modeling.
@@
In contrast to the encoder-decoder architecture adopted by Unified-IO and Parti, the decoder-only architecture leads to a simpler and more extensible approach for multimodal generative modeling. By transforming text and image inputs into a unified sequence of discrete tokens, the understanding and generation within and across modalities can be elegantly unified, forming a rosy direction towards general intelligence. During training, mGPT models the conditional probability $p(x_{t}|x_{1},...,x_{t-1})$ of multimodal sequences using the standard next-token prediction objective. Additionally, Chameleon applies z-loss (palm) to stabilize the training.
@@
Although Chameleon demonstrates potential for joint image and text understanding within one decoder-only transformer, its image generation ability remains inferior to state-of-the-art diffusion-based frameworks (sd3; pixart-sigma; hunyuandit; kolors; lumina-next) in both quality and resolution flexibility. Moreover, it is worth noting that the image generation ability is even absent in the open-source version of Chameleon. Additionally, the capabilities of Chameleon are confined to vision-language and text-only tasks, excluding a broader range of vision-centric tasks. These include classic visual recognition tasks such as segmentation and depth prediction, as well as creative visual generation tasks like controllable generation and image editing. Lumina-mGPT is built upon Chameleon to unlock its full potential for flexible photorealistic image generation and to become a versatile vision generalist.